An electrically neutral chlorine atom has 17 electrons orbiting the nucleus. One of the stable isotopes of this element has 18 neutrons in the nucleus. What is the atomic mass number of this isotope?
1
Expert's answer
2020-04-20T19:01:54-0400

M = p + n = e + n = 17 + 18 = 35 (Da)